# Stereographic Projections for Designs on the Sphere

### v1

@@ 4 Discussion
There are, in fact, many ways of representing an arrangement of points on the three-dimensional sphere in two dimensions, that is on paper, other than the stereogram. Thus, most simply, a drawing of the sphere with the points added or a sketch of the solid with vertices defined by the points can be used but these approaches are tedious to implement and not helpful in comparative studies. Alternatively, since a set of points on the sphere defines a convex polytope, the points can be represented by a 3-vertex-connected planar graph, that is a polyhedral graph, but the sense of the spatial arrangement is then lost. There are, in addition, projections, other than the stereographic projection, such as the orthographic and the gnomonic projections, which can be invoked but these distort the symmetry and, in the latter case, project the points onto an infinite plane tangential to the sphere. On balance therefore, none of the approaches discussed here offers the ease of use, the sense of the spherical space and the flexibility provided by the stereogram, as is required within the context of experimental design.
@@
Finally, it would be remiss not to mention the use of computer-generated three-dimensional models which represent points and solids with vertices on the sphere and which are available in, for example, the programming language Mathematica (2022). Such models can be rotated in space and thereby examined but it is nevertheless not easy to identify the inbuilt symmetry. The question also arises as to whether points on the four-dimensional sphere can be represented as three- or two-dimensional objects. There are ways of doing this, as for example with a Schlegel diagram, but these are complicated to interpret and not immediately useful in design.
Acknowledgements
@@
The author would like to thank the University of Cape Town and the National Research Foundation (NRF) of South Africa, grant (UID) 119122, for financial support. Any opinion, finding and conclusion or recommendation expressed in this material is that of the author and the NRF does not accept liability in this regard.

### v2

@@ 5. Discussion
The focus of the present study is on the use of the stereographic projection within the context of design in order to map points on the three-dimensional unit sphere onto the stereogram in two dimensions. The stereogram gives a clear visualisation of the spatial arrangement of the points on the sphere, as demonstrated here within the context of spherical -designs and regression modelling. There are, in fact, many ways of representing an arrangement of points on the three-dimensional sphere in two dimensions, that is on paper, other than the stereogram. Thus, most simply, a drawing of the sphere with the points added or a sketch of the solid with vertices defined by the points can be used but these approaches are tedious to implement and not helpful in comparative studies. Alternatively, since a set of points on the sphere defines a convex polytope, the points can be represented by a 3-vertex-connected planar graph, that is a polyhedral graph, but the sense of the spatial arrangement is then lost. There are, in addition, projections, other than the stereographic projection, such as the orthographic and the gnomonic projections, which can be invoked but these distort the symmetry and, in the latter case, projects the points onto an infinite plane tangential to the sphere. On balance therefore, none of the approaches discussed here offers the ease of use, the sense of the spherical space and the flexibility provided by the stereogram, as is required within the context of design.
@@
It would be remiss not to mention the use of computer-generated three-dimensional models which represent points and vertices of solids on the sphere and which are available in, for example, the website devoted to spherical -designs of Gräf (2011) and the programming language Mathematica (2024). Such models can be rotated in space and thereby examined but it is nevertheless not easy to identify the inbuilt symmetry. The question also arises as to whether points on the four-dimensional sphere can be represented as three- or two-dimensional objects. There are ways of doing this, as for example with a Schlegel diagram, but these are complicated to interpret and not immediately useful in design.
The applications presented in this study are of interest in themselves. Spherical -designs are researched extensively from a theoretical perspective in algebraic combinatorics and find application in technologies such as ambisonics. In addition, the designs are often cited, albeit somewhat peripherally, within the context of quantum theory. Designs for regression models with design space the unit ball are of interest in the area of design of experiments and, more specifically, in optimal design but tend to be represented, at best, as points on three-dimensional spheres. Stereograms can therefore be used to great advantage in research papers and, more particularly, in text books on experimental design to represent the spherical components of designs on the ball.
@@
Acknowledgements The author would like to thank the University of Cape Town and the National Research Foundation (NRF) of South Africa, grant (UID) 119122, for financial support. Any opinion, finding and conclusion or recommendation expressed in this material is that of the author and the NRF does not accept liability in this regard.